Suppose you just received a shipment of fifteen televisions. Four of the televisions are defective. If two televisions are randomly​ selected, compute the probability that both televisions work. What is the probability at least one of the two televisions does not​ work?
1
Expert's answer
2019-10-16T11:10:03-0400

In how many ways can 2 televisions be selected from 15?


"\binom{15}{2}={15! \over 2!(15-2)!}={15\cdot 14 \over 1\cdot 2}=105"

In how many ways can 0 defective televisions be selected from 4?


"\binom{4}{0}={4! \over 0!(4-0)!}=1"

In how many ways can 2 non-defective televisions be selected from 11?


"\binom{11}{2}={11! \over 2!(11-2)!}={11\cdot 10 \over 1\cdot 2}=55"

Compute the probability that both televisions work


"P(both\ work)={\binom{4}{0}\binom{11}{2} \over \binom{15}{2}}={1\cdot 55 \over 105}={11 \over 21}\approx0.5238"

What is the probability at least one of the two televisions does not​ work?


"P(at\ least \ 1\ does\ not\ work)=1-P(both\ work)=""=1-{11 \over 21}={10 \over 21}\approx0.4762"